# Implicit Safe Set Algorithm for Provably Safe Reinforcement Learning

### v1

@@ 5.1.2 Proof of Proposition 1
${\mathcal{U}_{S}^{D}}(x)$ :=$:=\{u\in\mathcal{U}\mid\max_{i}\phi_{i}(f(x,u))\leq\max_{i}\phi_{i}(x)\}$ (19a)
=$=\{u\in\mathcal{U}\mid\max_{i}\phi_{i}(f(x,u))\leq\phi_{j}(x)\}$ (19b)
$=\cap_{i}\{u\in\mathcal{U}\mid\phi_{i}(f(x,u))\leq\phi_{j}(x)\}$ (19c)
$={\mathcal{U}_{S}^{D}}_{j}(x)\neq\emptyset$ (19d)
@@
${\mathcal{U}_{S}^{D}}(x)$ :=$:=\{u\in\mathcal{U}\mid\max_{i}\phi_{i}(f(x,u))\leq 0\}$ (20a)
$=\cap_{i}\{u\in\mathcal{U}\mid\phi_{i}(f(x,u))\leq 0\}$ (20b)
$=\mathcal{U}\neq\emptyset$ (20c)
@@ 5.2 Feasibility of ISSA
Figure 3: Illustration of the grid sampling to find an anchor control point.
@@
$T^{a}<\prod\limits_{i=1}^{n_{u}}\lceil\frac{\zeta_{[i]}}{s^{*}_{l}}\rceil~{},$ (21)
@@
Figure 4: Illustration of the case when it is unable to find $u^{*}$.
@@
In this case, the second AdamBA process is evoked (line 14 of Algorithm 2). Since no solution is returned from the first AdamBA process (12 of Algorithm 2), where we start from $u^{r}$ and exponentially outreach along the direction $\vec{v}_{a}=\frac{u^{a}-u^{r}}{||u^{a}-u^{r}||}$, then all the searched control point along $\vec{v}_{a}$ is UNSAFE.
@@
$\begin{cases}u_{s}^{1}=u^{r}+\frac{\|u^{a}-u^{r}\|}{4}\vec{v}_{a}=\frac{3u^{r}}{4}+\frac{u^{a}}{4}\\ u_{s}^{2}=u^{r}+\frac{3\|u^{a}-u^{r}\|}{4}\vec{v}_{a}=\frac{u^{r}}{4}+\frac{3u^{a}}{4}\end{cases}$ (23)
@@
During the second AdamBA process (line 14 of Algorithm 2), we start from $u^{a}$ and exponentially outreach along the direction $\vec{v}_{r}=\frac{u^{r}-u^{a}}{||u^{r}-u^{a}||}$ with initial step size $\beta=\frac{\|u^{r}-u^{a}\|}{4}$. Hence, denote $\bar{u}_{s}^{1}$ as the first sample point along $\vec{v}_{r}$, and $\bar{u}_{s}^{1}$ satisfies
$\bar{u}_{s}^{1}=u^{a}+\frac{\|u^{r}-u^{a}\|}{4}\vec{v}_{r}=\frac{u^{r}}{4}+\frac{3u^{a}}{4},$ (24)
@@ 5.2.2 Proof of Proposition 2
5.3 Proof of Theorem 1
Before we prove Theorem 1, we first define set $\mathcal{X}_{S}^{D}:=\{x|\phi(x)\leq 0\}$, and we start with a preliminary result regarding $\mathcal{X}_{S}^{D}$ that is useful for proving the main theorem:
@@ 5.3 Proof of Theorem 1
If the control system satisfies Assumption 1 and Assumption 2, and the safety index design follows the rule described Section 4.1, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{X}_{S}^{D}$.
@@
If the control system satisfies the assumptions in Assumption 1 and Assumption 2, and the safety index design follows the rule described Section 4.1, then we can ensure the system has nonempty set of safe control at any state by Proposition 1. By Proposition 2, the implicit safe set algorithm can always find local optima solution of (5). The local optima solution always satisfies the constraint $\phi(f(x_{t},u_{t}))\leq\max\{\phi(x_{t})-\eta,0\}$, which indicates that 1) if $\phi(x_{t_{0}})\leq 0$, then $\phi(x_{t})\leq 0,\forall t\geq t_{0}$. Note that $\phi(x)\leq 0$ demonstrates that $x\in\mathcal{X}_{S}^{D}$. ∎
5.3.1 Proof the Theorem 1
@@
In this case, $\mathcal{S}=\mathcal{X}_{S}^{D}$. According to Lemma 6, If the control system satisfies the assumptions in Assumption 1 and Assumption 2, and the safety index design follows the rule described Section 4.1, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{X}_{S}^{D}$ and hence $\mathcal{S}$.
@@
Firstly, we consider the case where ${\phi_{0}}_{i}(x_{t})<0$. Note that ${\phi_{0}}_{i}(x_{t+1})={\phi_{0}}_{i}(x_{t})+\dot{\phi_{0}}_{i}(x_{t})dt+\frac{\ddot{\phi_{0}}_{i}(x_{t})dt^{2}}{2!}+\cdots$, since the state space and control space are both bounded, and $dt\rightarrow 0$ according to Assumption 1, we have ${\phi_{0}}_{i}(x_{t+1})\rightarrow{\phi_{0}}_{i}(x_{t})\leq 0$.
@@
In summary, by discussing the two cases of whether $\mathcal{X}_{S}^{D}$ is the subset of $\mathcal{X}_{S}$, we have proven that if the control system satisfies the assumptions in Assumption 1 and Assumption 2, and the safety index design follows the rule described in Section 4.1, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}\subseteq\mathcal{X}_{S}$.
6 Implicit Safe Set Algorithm for Systems with Non-Negligible Sampling Time
@@
$\begin{split}\phi(x_{t+1})=\phi(x_{t})+\dot{\phi}(x_{t})dt+O(dt^{2}),\dot{\phi}(x_{t})=\lim_{\tau\rightarrow 0^{+}}\frac{\phi(x_{t+\tau/dt})-\phi(x_{t})}{\tau}.\end{split}$ (26)
Therefore, the assumptions concerning sparse obstacle environment in Assumption 2 should also be justified. When $dt\rightarrow 0$, $\phi_{t+1}\rightarrow\phi_{t}$ according to (26), thus the safety-critical condition should be $\phi\geq 0$. However, when $dt$ is non-negligible, there will be a gap between $\phi_{t+1}$ and $\phi_{t}$ which can be represented as $\Delta_{\phi_{max}}$. Then the safety-critical condition should be $\phi+\Delta_{\phi_{max}}\geq 0$, where $\Delta_{\phi_{max}}$ is the maximum possible $\phi$ change magnitude at one time step, i.e. $\Delta_{\phi_{max}}=\max_{x,u,t}|\phi_{t+1}-\phi_{t}|$.
@@
  • Offline: design a safety index that ensures $\mathcal{U}_{S}^{D}(x)$ is nonempty for all $x$ for discrete-time system
@@
6.1 Safety Index Design Rule for Discrete-Time System
@@
  • safety index design rule from Section 4.1 sets $\eta=0$, hence robot is unable to converge to the safe set (finite time convergence) via solving (5) if the robot starts from the unsafe state, i.e. $\phi(x_{t_{0}})>0$.
  • safety index design rule from Section 4.1 relies negligible sampling time assumption to guarantee the non-emptiness of the set of safe control, which doesn’t hold in discrete-time system
@@ 6.1.1 Assumption
1) The discrete-time system time step satisfies the condition: $\frac{a_{min}}{2}+\frac{v_{max}}{4dt}>(a_{m}+v_{max}w_{m})(-\frac{a_{min}}{v_{max}}+w_{m})dt$, where $a_{m}=\max\{-a_{min},a_{max}\}$ and $w_{m}=\max\{-w_{min},w_{max}\}$; 2) The acceleration and angular velocity keep constant between two consecutive time steps, i.e. $a_{[t:t+1)}=a_{t}$. 3) At any given time, there can at most be one obstacle becoming safety-critical, such that $\phi\geq-\Delta_{\phi_{max}}$ where $\Delta_{\phi_{max}}$ is the maximum possible $\phi$ change magnitude at one time step, i.e. $\Delta_{\phi_{max}}=\max_{x,u,t}|\phi_{t+1}-\phi_{t}|$ (Sparse Obstacle Environment).
These assumptions are easy to meet in practice. The first assumption ensures a bounded sampling time for discrete-time systems, which will be leveraged for the theoretical proofs in Section 7. The second assumption is the zero-order hold (ZOH) for discrete-time control signal. The third assumption enables safety index design rule applicable with multiple moving obstacles in discrete-time system.
6.1.2 Safety Index Design Rule
Following the rules in (?), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$. Our goal is to choose these parameters such that $\mathcal{U}_{S}^{D}(x)$ is always nonempty for all possible states in discrete-time system. Although similar to the safety index design rule in continuous-time systems, the safety index design rule for discrete-time systems is updated mainly from three aspects.
  • We use nontrivial positive $\sigma$ as an safety margin to enable the forward invariance to safe set $\mathcal{S}$ for discrete-time systems.
@@ 6.1.2 Safety Index Design Rule
  • We use nontrivial positive $\eta$ to enable the finite time convergence to safe set $\mathcal{S}$ when discrete-time systems start from unsafe states.
@@
$\sigma$ $>-\dot{d}^{*}_{min}dt$ (27a)
$\frac{\frac{\eta_{0}}{dt}+v_{\max}}{k}$ $\leq\min\{-a_{min},a_{max}\}~{},$ (27b)
@@
6.2 Implicit Safe Set Algorithm for Discrete-Time System
The ultimate goal of implicit safe set algorithm for discrete-time systems is to ensure (i) forward invariance to $\mathcal{X}_{S}^{D}$ when systems start from safe states, and (ii) finite time convergence to $\mathcal{X}_{S}^{D}$ when systems start from unsafe states. The discrete-time system implicit safe set algorithm starts by applying ISSA (Algorithm 2), then applies an additional convergence trigger (CTrigger) algorithm to filter ISSA solution when $|\cos(\alpha)|$ is small and $\phi(x)>0$.
The main body of CTrigger algorithm is summarized in Algorithm 3. The inputs for CTrigger algorithm include (i) system state ($x$), (ii) the safe control from ISSA ($u$), (iii) the angle between the robot’s heading vector and the vector from the robot to the obstacle ($\alpha$), (iv) the minimum relative acceleration ($a_{min}$), (v) the maximum relative acceleration ($a_{max}$), (vi) ($w_{trigger}$), where $w_{trigger}=\inf_{v\geq\frac{v_{max}}{2}}\sup_{u}|w|$, and (vii) ($\Delta_{\min}$), where
$\Delta_{\min}=\inf_{|\cos(\alpha)|\leq\frac{\sqrt{3}}{2},|w|\geq\frac{|w_{trigger}|}{2}}|\Delta\cos(\alpha)|$, and $\Delta\cos(\alpha)$ denotes the change of $\cos(\alpha)$ between two consecutive time steps. Note that $a_{min},a_{max},w_{trigger}$ and $\Delta_{\min}$ are fundamental system properties, which can be evaluated offline.
@@ 6.2 Implicit Safe Set Algorithm for Discrete-Time System
The core idea of CTrigger algorithm is to enable finite time convergence via preventing $|\cos(\alpha)|$ from approaching zero (according to (28)). Intuitively this is preventing the robot from moving in a circle around the obstacle at a constant distance, and thus preventing non-convergence. Specifically, CTrigger filters the safe control solution from ISSA when $|\cos(\alpha)|$ less than a threshold, i.e. $\min\{\frac{\sqrt{3}}{2},\frac{\Delta_{\min}}{2}\}$, which is accomplished via two main steps. I. Generate safe control with non-trivial acceleration to enable the system to gain enough speed (Line 4 of Algorithm 3); II. Generate safe control with non-trivial angular velocity to push $|\cos(\alpha)|$ away from zero (Line 6 of Algorithm 3). And we show in Theorem 2 that ISSA with CTrigger algorithm ensure forward invariance and finite time convergence to the set $\mathcal{S}:=\mathcal{X}_{S}\cap\mathcal{X}_{S}^{D}$ for discrete-time systems.
@@
Note that there are three heuristic based hyper-parameters in Algorithm 3, including $\frac{\Delta_{min}}{2}$, $\frac{\min\{-a_{min},a_{max}\}}{2}$, and $\frac{|w_{trigger}|}{2}$, and the selection heuristics are summarized as following: 1) $\frac{\Delta_{min}}{2}$ ensures that the triggering control will only be activated when $|\cos(\alpha)|$ is sufficiently small ; 2) $\frac{\min\{-a_{min},a_{max}\}}{2}$ ensures that the candidate safe triggering control set for acceleration (Line 4 in Algorithm 3) is non-trivial; 3) $\frac{|w_{trigger}|}{2}$ ensures that the candidate safe triggering control set for rotation (Line 6 in Algorithm 3) is non-trivial. Although $a_{min},a_{max},w_{trigger}$ and $\Delta_{\min}$ are the fundamental system properties, it is possible to change the denominators of the hyper-parameters, e.g. $\frac{|w_{trigger}|}{3}$, or $\frac{\min\{-a_{min},a_{max}\}}{5}$. However, there are a trade-offs when we change the denominators of the hyper-parameters. Specifically, system will get out of Singularity State faster if the denominator decreases, whereas it may take longer sampling time for finding a valid triggering control. In practice, we find setting denominator as $2$ works well for balancing the aforementioned trade-offs.
7 Theoretical Results for ISSA When the Sampling Time is Non-Negligible

### v2

@@ 5.1.2. Proof of Proposition 2
(19a) ${\mathcal{U}_{S}^{D}}(x)$ $:=\{u\in\mathcal{U}\mid\max_{i}\phi_{i}(f(x,u))\leq\max_{i}\phi_{i}(x)\}$
(19b) $=\{u\in\mathcal{U}\mid\max_{i}\phi_{i}(f(x,u))\leq\phi_{j}(x)\}$
(19c) $=\cap_{i}\{u\in\mathcal{U}\mid\phi_{i}(f(x,u))\leq\phi_{j}(x)\}$
(19d) $={\mathcal{U}_{S}^{D}}_{j}(x)\neq\emptyset$
@@
(20a) ${\mathcal{U}_{S}^{D}}(x)$ $:=\{u\in\mathcal{U}\mid\max_{i}\phi_{i}(f(x,u))\leq 0\}$
(20b) $=\cap_{i}\{u\in\mathcal{U}\mid\phi_{i}(f(x,u))\leq 0\}$
(20c) $=\mathcal{U}\neq\emptyset$
@@ 5.2. Feasibility of ISSA
Figure 4. Illustration of the grid sampling to find an anchor control point.
@@
(21) $T^{a}<\prod\limits_{i=1}^{n_{u}}\lceil\frac{\zeta_{[i]}}{s^{*}_{l}}\rceil~,$
@@
Figure 5. Illustration of the case when it is unable to find $u^{*}$.
@@
[…]
@@
(23) $\begin{cases}u_{s}^{1}=u^{r}+\frac{\|u^{a}-u^{r}\|}{4}\vec{v}_{a}=\frac{3u^{r}}{4}+\frac{u^{a}}{4}\\ u_{s}^{2}=u^{r}+\frac{3\|u^{a}-u^{r}\|}{4}\vec{v}_{a}=\frac{u^{r}}{4}+\frac{3u^{a}}{4}\end{cases}$
@@
[…]
(24) $\bar{u}_{s}^{1}=u^{a}+\frac{\|u^{r}-u^{a}\|}{4}\vec{v}_{r}=\frac{u^{r}}{4}+\frac{3u^{a}}{4},$
@@ 5.2.2. Proof of Proposition 5
5.3. Proof of Theorem 1
Before we prove Theorem 1, we start with a preliminary result regarding $\mathcal{X}_{S}^{D}$ that is useful for proving the main theorem:
@@ 5.3. Proof of Theorem 1
If the control system satisfies 1 and 2, and the safety index design follows the rule described Section 4.1, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{X}_{S}^{D}$.
@@
If the control system satisfies the assumptions in 1 and 2, and the safety index design follows the rule described Section 4.1, then we can ensure the system has a nonempty set of safe control at any state by Proposition 2. By Proposition 5, the implicit safe set algorithm can always find a local optima solution of (5). The local optima solution always satisfies the constraint $\phi(f(x_{t},u_{t}))\leq\max\{\phi(x_{t})-\eta,0\}$, which indicates that 1) if $\phi(x_{t_{0}})\leq 0$, then $\phi(x_{t})\leq 0,\forall t\geq t_{0}$. Note that $\phi(x)\leq 0$ demonstrates that $x\in\mathcal{X}_{S}^{D}$. ∎
Now the proof the Theorem 1 is detailed below.
@@
In this case, $\mathcal{S}=\mathcal{X}_{S}^{D}$. According to Lemma 9, If the control system satisfies the assumptions in 1 and 2, and the safety index design follows the rule described Section 4.1, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{X}_{S}^{D}$ and hence $\mathcal{S}$.
@@
Firstly, we consider the case where ${\phi_{0}}_{i}(x_{t})<0$. Note that ${\phi_{0}}_{i}(x_{t+1})={\phi_{0}}_{i}(x_{t})+\dot{\phi_{0}}_{i}(x_{t})dt+\frac{\ddot{\phi_{0}}_{i}(x_{t})dt^{2}}{2!}+\cdots$, since the state space and control space are both bounded, and $dt\rightarrow 0$ according to 1, we have ${\phi_{0}}_{i}(x_{t+1})\rightarrow{\phi_{0}}_{i}(x_{t})\leq 0$.
@@
In summary, by discussing the two cases of whether $\mathcal{X}_{S}^{D}$ is the subset of $\mathcal{X}_{S}$, we have proven that if the control system satisfies the assumptions in 1 and 2, and the safety index design follows the rule described in Section 4.1, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}\subseteq\mathcal{X}_{S}$.
6. Implicit Safe Set Algorithm for Systems with Non-Negligible Sampling Time
@@
(26) $\begin{split}\phi(x_{t+1})=\phi(x_{t})+\dot{\phi}(x_{t})dt+O(dt^{2}),\dot{\phi}(x_{t})=\lim_{\tau\rightarrow 0^{+}}\frac{\phi(x_{t+\tau/dt})-\phi(x_{t})}{\tau}.\end{split}$
Therefore, the assumptions concerning sparse obstacle environment in 2 should also be justified. When $dt\rightarrow 0$, $\phi_{t+1}\rightarrow\phi_{t}$ according to (26), thus the safety-critical condition should be $\phi\geq 0$. However, when $dt$ is non-negligible, there will be a gap between $\phi_{t+1}$ and $\phi_{t}$ which can be represented as $\Delta_{\phi_{max}}$. Then the safety-critical condition should be $\phi+\Delta_{\phi_{max}}\geq 0$, where $\Delta_{\phi_{max}}$ is the maximum possible $\phi$ change magnitude at one time step, i.e. $\Delta_{\phi_{max}}=\max_{x,u,t}|\phi_{t+1}-\phi_{t}|$.
@@
  • Offline: design a safety index that ensures $\mathcal{U}_{S}^{D}(x)$ is nonempty for all $x$ for discrete-time system.
@@
6.1. Safety Index Design Rule for Discrete-Time System
@@
  • the safety index design rule from Section 4.1 sets $\eta=0$, hence the robot is unable to converge to the safe set (finite time convergence) via solving (5) if the robot starts from the unsafe state, i.e. $\phi(x_{t_{0}})>0$.
  • the safety index design rule from Section 4.1 relies on the negligible sampling time assumption to guarantee the non-emptiness of the set of safe control, which doesn’t hold in discrete-time system.
@@ 6.1.1. Assumption
1) The discrete-time system time step satisfies the condition: $\frac{a_{min}}{2}+\frac{v_{max}}{4dt}>(a_{m}+v_{max}w_{m})(-\frac{a_{min}}{v_{max}}+w_{m})dt$, where $a_{m}=\max\{-a_{min},a_{max}\}$ and $w_{m}=\max\{-w_{min},w_{max}\}$; 2) The acceleration and angular velocity keep constant between two consecutive time steps, i.e. $a_{[t:t+1)}=a_{t}$. 3) At any given time, there can at most be one obstacle becoming safety-critical, such that $\phi\geq-\Delta_{\phi_{max}}$ where $\Delta_{\phi_{max}}=\max_{x,u,t}|\phi_{t+1}-\phi_{t}|$ is the maximum possible $\phi$ change magnitude at one time step (Sparse Obstacle Environment).
These assumptions are easy to meet in practice. The first assumption ensures a bounded sampling time for discrete-time systems, which will be leveraged for the theoretical proofs in Section 7. The second assumption assumes zero-order hold (ZOH) for discrete-time control signals. The third assumption enables safety index design rule applicable with multiple moving obstacles in discrete-time system. $\Delta_{\phi_{max}}$ is bounded mainly because the system dynamics are bounded and the sampling time is also bounded.
6.1.2. Safety Index Design Rule
Following the rules in (liu2014control, 25), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$. Our goal is to choose these parameters such that $\mathcal{U}_{S}^{D}(x)$ is always nonempty for all possible states in the discrete-time system. Although similar to the safety index design rule in continuous-time systems, the safety index design rule for discrete-time systems is updated mainly from three aspects.
  • We use nontrivial positive $\sigma$ as an safety margin to enable the forward invariance to the invariantly safe set $\mathcal{S}$ for discrete-time systems.
@@ 6.1.2. Safety Index Design Rule
  • We use nontrivial positive $\eta$ to enable the finite time convergence to the invariantly safe set $\mathcal{S}$ when the discrete-time system starts from initially unsafe states or initially safe but inevitably unsafe states in $\mathcal{X}_{S}$.
@@
(27a) $\sigma$ $>-\dot{d}^{*}_{min}dt$
(27b) $\frac{\frac{\eta_{0}}{dt}+v_{\max}}{k}$ $\leq\min\{-a_{min},a_{max}\}~,$
@@
6.2. Implicit Safe Set Algorithm for Discrete-Time System
The ultimate goal of the implicit safe set algorithm for discrete-time systems is to ensure (i) forward invariance in $\mathcal{S}=\mathcal{X}_{S}\cap\mathcal{X}_{S}^{D}$ when systems start from safe states, and (ii) finite time convergence to $\mathcal{S}$ when systems start from unsafe states. The discrete-time implicit safe set algorithm starts by applying ISSA (Algorithm 2), then applies an additional convergence trigger (CTrigger) algorithm to filter ISSA solution when $|\cos(\alpha)|$ is small and $\phi(x)>0$.
The main body of CTrigger algorithm is summarized in Algorithm 3. The inputs for CTrigger algorithm include:
  1. (1)
    system state ($x$),
  2. (2)
    the safe control from ISSA ($u$),
  3. (3)
    the angle between the robot’s heading vector and the vector from the robot to the obstacle ($\alpha$),
  4. (4)
    the minimum relative acceleration ($a_{min}$),
  5. (5)
    the maximum relative acceleration ($a_{max}$),
  6. (6)
    the triggering angular velocity ($w_{trigger}$), where $w_{trigger}=\inf_{v\geq\frac{v_{max}}{2}}\sup_{u}|w|$, and
  7. (7)
    the triggering angle ($\Delta_{\min}$), where $\Delta_{\min}=\inf_{|\cos(\alpha)|\leq\frac{\sqrt{3}}{2},|w|\geq\frac{|w_{trigger}|}{2}}|\Delta\cos(\alpha)|$, and $\Delta\cos(\alpha)$ denotes the change magnitude of $\cos(\alpha)$ between two consecutive time steps.
Note that $a_{min},a_{max},w_{trigger}$ and $\Delta_{\min}$ are fundamental system properties, which can be evaluated offline.
@@ 6.2. Implicit Safe Set Algorithm for Discrete-Time System
The core idea of CTrigger algorithm is to enable finite time convergence via preventing $|\cos(\alpha)|$ from approaching zero (according to (28)). Intuitively, this is preventing the robot from moving in a circle around the obstacle at a constant distance, and thus preventing non-convergence. Specifically, CTrigger filters the safe control solution from ISSA when $|\cos(\alpha)|$ is less than a threshold, i.e. $\min\{\frac{\sqrt{3}}{2},\frac{\Delta_{\min}}{2}\}$, which is accomplished via two main steps. I. Generate safe control with non-trivial acceleration to enable the system to gain enough speed (Line 4 of Algorithm 3); II. Generate safe control with non-trivial angular velocity to push $|\cos(\alpha)|$ away from zero (Line 6 of Algorithm 3). And we show in Theorem 1 that ISSA with CTrigger algorithm ensures forward invariance and finite time convergence to the set $\mathcal{S}=\mathcal{X}_{S}\cap\mathcal{X}_{S}^{D}$ for discrete-time systems.
@@
Note that there are three heuristic based hyper-parameters in Algorithm 3, including $\frac{\Delta_{min}}{2}$, $\frac{\min\{-a_{min},a_{max}\}}{2}$, and $\frac{|w_{trigger}|}{2}$, and the selection heuristics are summarized as following: 1) $\frac{\Delta_{min}}{2}$ ensures that the triggering control will only be activated when $|\cos(\alpha)|$ is sufficiently small ; 2) $\frac{\min\{-a_{min},a_{max}\}}{2}$ ensures that the candidate safe triggering control set for acceleration (Line 4 in Algorithm 3) is non-trivial; 3) $\frac{|w_{trigger}|}{2}$ ensures that the candidate safe triggering control set for rotation (Line 6 in Algorithm 3) is non-trivial. Although $a_{min},a_{max},w_{trigger}$ and $\Delta_{\min}$ are the fundamental system properties, it is possible to change the denominators of the hyper-parameters, e.g. $\frac{|w_{trigger}|}{3}$, or $\frac{\min\{-a_{min},a_{max}\}}{5}$. However, there are trade-offs when we change the denominators of the hyper-parameters. Specifically, system will get out of Singularity State faster if the denominator decreases, whereas it may take longer sampling time for finding a valid triggering control. In practice, we find setting the denominator as $2$ works well for balancing the aforementioned trade-offs.
7. Theoretical Results for ISSA When the Sampling Time is Non-Negligible